OpenStudy (anonymous):

What is the fourth term of (x + y)8?

mathslover (mathslover):

just note the basic formula that is BINOMIAL THEOREM :

OpenStudy (anonymous):

It's the 8th row. \[\left(\begin{matrix}8 \\ 4\end{matrix}\right) = \frac{8!}{4!4!}\]Solve this for the coefficient. Then, following the binomial theorem like @mathslover said, place an x^5 and y^3 in front to get the whole term. \[\frac{8!}{4!4!} \times x^{5}\times y^{3} = ?\]

OpenStudy (anonymous):

Or if you have a lot of paper, use Pascal's triangle!

OpenStudy (anonymous):

56x4y4??

OpenStudy (anonymous):

@sterna 56x4y4??

OpenStudy (anonymous):

i mean 56x5y3

OpenStudy (anonymous):

8! means 8*7*6*5*4*3*2*1, but you probably know that. If you divide it by 4!, you just cancel the 4*3*2*1 and you're left with 8*7*6*5. I think mathcalc is saying to divide by 4! again. Try that : divide (8*7*6*5) by (4*3*2*1) and see what you get.

OpenStudy (turingtest):

\[(x+y)^n=\sum_{k=0}^n\binom nk x^{n-k}y^k\]where \[\binom nk={n!\over k!(n-k)!}\]the \(k+1^{th}\) term is given by \[\binom nk x^{n-k}y^k\]

OpenStudy (turingtest):

you want the \(4^{th}\) term, so \(k=3\)

OpenStudy (turingtest):

...and \(n=8\)

OpenStudy (anonymous):

However, I get something different using the trusty old Pascal's triangle.

OpenStudy (turingtest):

\[\binom 83x^{8-3}y^3={8!\over3!(8-3)!}x^5y^3={8\cdot7\cdot\cancel6\over\cancel{3\cdot2}}x^5y^3=56x^5y^3\]hooray, math works!

OpenStudy (turingtest):

so both of the above are wrong :P

OpenStudy (anonymous):

Nicely done: that's what Pascal told me as well - so does that help 4everJames?

OpenStudy (anonymous):

Oops. Sorry. I knew I forgot something :/

OpenStudy (anonymous):

@TuringTest Thanks for clarifying my thoughts :)